OpenStudy (anonymous):

Rearrange the following to make y the subject t = √3y 2 – 5

OpenStudy (anonymous):

\[\sqrt{3y} * 2 - 5\]?

OpenStudy (whpalmer4):

@joshhhunt if you're just going to use a square root character like that, at least put parentheses around the stuff under the square root sign so there is no question about what the problem is...

OpenStudy (whpalmer4):

Better yet, click on the Equation button and format it nicely...make it easier to help you.

OpenStudy (anonymous):

Hi, I'm trying to learn how to get the answer. My bad on not making it clear: t = \[\sqrt{3y ^{2}-5}\] (Im new)

OpenStudy (whpalmer4):

Not a problem, just trying to help you help us help you :-)

OpenStudy (whpalmer4):

so you have \[t=\sqrt{3y^2-5}\] and you want to solve for \(y\) in terms of \(t\), correct?

OpenStudy (anonymous):

The task is to rearrange it so that y is the subject if that helps.

OpenStudy (whpalmer4):

I would square both sides of the equation. Remember that \((\sqrt{x})^2=x\)

OpenStudy (whpalmer4):

What do you get after squaring both sides?

OpenStudy (anonymous):

I should know this but how do I go about that.. the previous questions like k = 2y + 15 – 4y I didn't need to square either sides to get the answer

OpenStudy (anonymous):

I understand how to do them but the square root throws me off for this question

OpenStudy (whpalmer4):

If we square both sides, we have this: \[t^2 = (\sqrt{3y^2-5})^2\]but as I said, if you square the square root, when the dust settles, you just have what was under the square root sign: \[t^2=3y^2-5\]

OpenStudy (anonymous):

So I assume the next step would be to add 5 to both sides, right?

OpenStudy (whpalmer4):

That's what I would do.

OpenStudy (anonymous):

Ok so now I have \[t{^2} + 5=3y{^2}\] Then I would divide both sides by 3 and end up with this \[7t \div 3=y2\] Am I on the right track?

OpenStudy (anonymous):

Wait I've done something wrong, hold up.

OpenStudy (anonymous):

\[t{^2}+5 \div 3=y2\] ?

OpenStudy (whpalmer4):

\[t^2+5 = 3y^2\]\[\frac{t^2+5}{3} = y^2\]

OpenStudy (whpalmer4):

Or if you prefer, \[\frac{t^2}{3} + \frac{5}{3} = y^2\]

OpenStudy (whpalmer4):

But neither of those is equivalent to \[t^2+5/3 = y^2\]because division has a higher priority than addition. Remember PEMDAS?

OpenStudy (anonymous):

Ahh wrong symbol, so what do I do with the 2 so that I am left with just y? \[y{^2}\]

OpenStudy (whpalmer4):

Now you take the square root of both sides: \[\sqrt{\frac{t^2+5}{3}} = y\]And then switch sides \[y=\sqrt{\frac{t^2+5}{3}}\] I assume that's what your instructor wants when saying "make y the subject"

OpenStudy (anonymous):

That looks right to me, thank you very much for your help.